# Convergence Rate of Particle System for Second-order PDEs On Wasserstein Space

### v1

@@ 2.2. Preliminary results
$\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})\leq C\left(\epsilon+\alpha(N)+\sqrt{\epsilon\alpha(N)}\right),$
@@
According to [33, Lemma 5.1,5.4], we have
@@
$\frac{1}{2}\partial_{\nu}\rho_{*}^{2}(\mu,\nu)(x)\leq\rho_{*}(\mu,\nu)\sqrt{\sum_{l\in{\mathbb{Z}}^{d}}|l|^{2}(1+|l|^{2})^{-k_{*}}},$
where the constant $\sqrt{\sum_{l\in{\mathbb{Z}}^{d}}|l|^{2}(1+|l|^{2})^{-k_{*}}}$ is finite due to our choice of $k_{*}$. Then according to Lemma 3.1 and Lemma 3.2,
$\left|D_{\mu}\Phi(t_{0},z_{0},\mu_{0})(x)\right|\leq\frac{\sqrt{\sum_{l\in{\mathbb{Z}}^{d}}|l|^{2}(1+|l|^{2})^{-k_{*}}}\left(\epsilon+\alpha(N)+\sqrt{\epsilon\alpha(N)}\right)}{\epsilon N}=:R(\epsilon,N).$
@@
$E(\epsilon,N)\leq C_{R(\epsilon,N)}\epsilon+\frac{C(R_{\epsilon,N})^{2}}{2\epsilon N}\rho^{2}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+C_{R(\epsilon,N)}\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+\frac{1}{2\epsilon N}\rho^{2}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0}),$
@@
Let us denote the terms on the right hand side of (3) by $(I),(II),(III),(IV)$ from the top to the bottom, and we have the following estimate.
@@
$D_{\mu}\Phi(\theta_{0})(\cdot)=\frac{1}{2\epsilon N}D_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(\cdot),$
According to [33, Lemma 5.4] the term on the right is in the Sobolev space $\mathbb{H}^{k_{*}-1}$. Indeed, according to (3.7)
$\lVert D_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(\cdot)\rVert^{2}_{H^{k_{*}-1}}=4\sum_{l\in{\mathbb{Z}}^{d}}(1+|l|^{2})^{k_{*}-1}|l|^{2}(1+|l|^{2})^{-2k_{*}}|F_{l}(\nu-\mu)|^{2}\leq 4\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0}),$
@@
$\lVert D_{\mu}\Phi(\theta_{0})\rVert_{\mathbb{H}^{k_{*}-1}}\leq\frac{1}{\epsilon N}\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0}).$
@@
$\lVert f\rVert_{\mathbb{H}^{k_{*}-1}}=\lVert g\circ(I_{d},h)\rVert_{\mathbb{H}^{k_{*}-1}}\leq C(R_{\epsilon,N})\lVert D_{\mu}\Phi(\theta_{0})\rVert_{\mathbb{H}^{k_{*}-1}},$
@@
$(II)$ $\leq\frac{C}{\epsilon N}\rho_{1-k_{*}}(\mu^{\boldsymbol{x}_{0}},\mu_{0})\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})$
$\leq\frac{1}{2\epsilon N}\rho^{2}_{1-k_{*}}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+\frac{C^{2}}{2\epsilon N}\rho^{2}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0}).$
@@
$(IV)=$ $-\frac{1}{2\epsilon N}\sum_{l\in{\mathbb{Z}}^{d}}\frac{|l|^{2}}{(1+|l|^{2})^{k_{*}}}F_{l}(\mu_{0}-\mu^{\boldsymbol{x}_{0}})\int e^{*}_{l}(x)\,(\mu_{0}-\mu^{\boldsymbol{x}_{0}})(dx)$
$=$ $-\frac{1}{2\epsilon N}\sum_{l\in{\mathbb{Z}}^{d}}\frac{|l|^{2}}{(1+|l|^{2})^{k_{*}}}|F_{l}(\mu_{0}-\mu^{\boldsymbol{x}_{0}})|^{2}$
$=$ $-\frac{1}{2\epsilon N}\rho^{2}_{1-k_{*}}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+\frac{1}{2\epsilon N}\rho^{2}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0}).$
Summing up the estimates for $(I),(II),(III),(IV)$, we conclude the result. ∎
@@
Together with (3.2), $\overline{V}^{N,\epsilon}(t,z,\mu)+(T-t)E(\epsilon,N)+C(\epsilon^{1/3}+\epsilon+\alpha(N)+\sqrt{\epsilon\alpha(N)})$ is a viscosity supersolution to (1.1). Then due to the comparison principle in Proposition 2.2, we have that
@@
Recalling that by the definition of inf-convolution, $V(t,z,\boldsymbol{x})\geq\overline{V}^{N,\epsilon}(t,z,\mu^{\boldsymbol{x}})$, and hence
@@
Choosing $\epsilon=\alpha(N)^{3}$, it can be easily checked that $R(\epsilon,N)$ is bounded independent of $N$, and therefore
$TE(\epsilon,N)+C(\epsilon^{1/3}+\epsilon+\alpha(N)+\sqrt{\epsilon\alpha(N)})$ $\leq C\alpha(N).$

### v2

@@ 2.2. Preliminary results
$\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})\leq C\left(\epsilon+\alpha(N)+\sqrt{\epsilon\alpha(N)}\right),$
@@
According to [31, Lemma 5.1,5.4], we have
@@
$\frac{1}{2}|D_{\nu}\rho_{*}^{2}(\mu,\nu)(x)|$ $\leq\rho_{*}(\mu,\nu)\sqrt{\sum_{l\in{\mathbb{Z}}^{d}}|l|^{2}(1+|l|^{2})^{-k_{*}}},$
$\frac{1}{2}|D^{2}_{\nu}\rho_{*}^{2}(\mu,\nu)(x,y)|$ $\leq\rho_{*}(\mu,\nu)\sqrt{\sum_{l\in{\mathbb{Z}}^{d}}|l|^{4}(1+|l|^{2})^{-k_{*}}}.$ (3.8)
where the constants $\sqrt{\sum_{l\in{\mathbb{Z}}^{d}}|l|^{2}(1+|l|^{2})^{-k_{*}}}$ and $\sqrt{\sum_{l\in{\mathbb{Z}}^{d}}|l|^{4}(1+|l|^{2})^{-k_{*}}}$ are finite due to our choice of $k_{*}$. Then according to Lemma 3.1 and Lemma 3.2,
$\left|D_{\mu}\Phi(t_{0},z_{0},\mu_{0})(x)\right|\leq\frac{\sqrt{\sum_{l\in{\mathbb{Z}}^{d}}|l|^{2}(1+|l|^{2})^{-k_{*}}}\left(\epsilon+\alpha(N)+\sqrt{\epsilon\alpha(N)}\right)}{\epsilon}=:R(\epsilon,N).$
@@
$E(\epsilon,N)\leq$ $C_{R(\epsilon,N)}\epsilon+C_{R(\epsilon,N)}(1+1/(\epsilon N))\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+\frac{1}{2\epsilon}\rho^{2}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})$
$+\frac{\epsilon C_{R(\epsilon,N)}^{2}}{4}\left(\frac{\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})}{\epsilon}+\frac{\rho^{k_{*}-1}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})}{\epsilon^{k_{*}-1}}\right)^{2},$
@@
Let us denote the terms on the right-hand side of (3) by $(I),(II),(III),$ and $(IV)$ from the top to the bottom, and we have the following estimate.
@@
$D_{\mu}\Phi(\theta_{0})(\cdot)=\frac{1}{2\epsilon}D_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(\cdot),$
According to [31, Lemma 5.4] the term on the right is in the Sobolev space $\mathbb{H}^{k_{*}-1}$. Indeed, according to (3.7)
$\lVert D_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(\cdot)\rVert^{2}_{k_{*}-1}$ $=4\sum_{l\in{\mathbb{Z}}^{d}}(1+|l|^{2})^{k_{*}-1}|l|^{2}(1+|l|^{2})^{-2k_{*}}|F_{l}(\nu-\mu)|^{2}$
$\leq 4\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0}),$
@@
$\lVert D_{\mu}\Phi(\theta_{0})\rVert_{{k_{*}-1}}\leq\frac{1}{\epsilon}\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0}).$
@@
$\lVert f\rVert_{{k_{*}-1}}=\lVert g\circ(I_{d},h)\rVert_{{k_{*}-1}}$ $\leq C(R_{\epsilon,N})\left(\lVert D_{\mu}\Phi(\theta_{0})\rVert_{{k_{*}-1}}+\lVert D_{\mu}\Phi(\theta_{0})\rVert_{{k_{*}-1}}^{k_{*}-1}\right)$
$\leq C(R_{\epsilon,N})\left(\frac{\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})}{\epsilon}+\frac{\rho^{k_{*}-1}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})}{\epsilon^{k_{*}-1}}\right)$
@@
$(II)$ $\leq C_{R(\epsilon,N)}\rho_{1-k_{*}}(\mu^{\boldsymbol{x}_{0}},\mu_{0})\left(\frac{\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})}{\epsilon}+\frac{\rho^{k_{*}-1}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})}{\epsilon^{k_{*}-1}}\right)$
$\leq\frac{1}{\epsilon}\rho^{2}_{1-k_{*}}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+\frac{\epsilon C_{R(\epsilon,N)}^{2}}{4}\left(\frac{\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})}{\epsilon}+\frac{\rho^{k_{*}-1}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})}{\epsilon^{k_{*}-1}}\right)^{2}.$
@@
$(IV)=$ $-\frac{1}{\epsilon}\sum_{l\in{\mathbb{Z}}^{d}}\frac{|l|^{2}}{(1+|l|^{2})^{k_{*}}}F_{l}(\mu_{0}-\mu^{\boldsymbol{x}_{0}})\int e^{*}_{l}(x)\,(\mu_{0}-\mu^{\boldsymbol{x}_{0}})(dx)$
$-\sum_{i=1}^{N}\frac{1}{2\epsilon N^{2}}D^{2}_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}_{0},x^{i}_{0})$
$=$ $-\frac{1}{\epsilon}\sum_{l\in{\mathbb{Z}}^{d}}\frac{|l|^{2}}{(1+|l|^{2})^{k_{*}}}|F_{l}(\mu_{0}-\mu^{\boldsymbol{x}_{0}})|^{2}-\sum_{i=1}^{N}\frac{1}{2\epsilon N^{2}}D^{2}_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}_{0},x^{i}_{0})$
$=$ $-\frac{1}{\epsilon}\rho^{2}_{1-k_{*}}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+\frac{1}{\epsilon}\rho^{2}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})-\sum_{i=1}^{N}\frac{1}{2\epsilon N^{2}}D^{2}_{\nu}\rho_{*}^{2}(\mu^{\boldsymbol{x}_{0}},\mu_{0})(x^{i}_{0},x^{i}_{0})$
$\leq$ $-\frac{1}{\epsilon}\rho^{2}_{1-k_{*}}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+\frac{1}{\epsilon}\rho^{2}_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0})+\frac{C}{\epsilon N}\rho_{*}(\mu^{\boldsymbol{x}_{0}},\mu_{0}),$
where we apply the estimate (3) to get the last inequality. Summing up the estimates for $(I),(II),(III),(IV)$, we conclude the result. ∎
@@
Together with (3), $\overline{V}^{N,\epsilon}(t,z,\mu)+(T-t)E(\epsilon,N)+C(\epsilon^{1/3}+\epsilon+\alpha(N)+\sqrt{\epsilon\alpha(N)})$ is a viscosity supersolution to (1.1). Then due to the comparison principle in Proposition 2.2, we have that
@@
Recalling that by the definition of inf-convolution, $V^{N}(t,z,\boldsymbol{x})\geq\overline{V}^{N,\epsilon}(t,z,\mu^{\boldsymbol{x}})$, and hence
@@
Choosing $\epsilon=\alpha(N)$, it can be easily checked that $R(\epsilon,N)$ is bounded independent of $N$, and therefore
$TE(\epsilon,N)+C(\epsilon^{1/3}+\epsilon+\alpha(N)+\sqrt{\epsilon\alpha(N)})$ $\leq C\alpha^{1/3}(N).$
@@
Acknowledgments
We would like to thank Joe Jackson for his comments on the first version of the paper.